# Structural evolution of quiescent galaxies at intermediate redshift at UV and red rest-frame wavelengths

### v1

@@ 5 Structural evolution of quiescent galaxies in two rest-frame wavelengths
Scaling relations for quiescent galaxies constrain the processes that drive their evolution. We first discuss bias that the observations limited to a single photometric band may introduce in the measured structural parameters of quiescent systems at $0.60.2<z<$ 0.6 (Sect. 5.1). We then use rest-frame estimates to explore two scaling relations (size-stellar mass and the Sérsic index-stellar mass) in two wavelength ranges (UV and red). We trace the corresponding structural evolution in size and central concentration of quiescent galaxies over the $0.60.2<z<$ 0.6 redshift interval (Sec. 5.2).
@@ 5.1 Sérsic profile parameters in observed and rest-frame wavelengths
Sizes decrease with wavelength and Sérsic index increase with wavelength for quiescent galaxies with $0.60.2<z<$ 0.6 (Sect. 4.1, Figures 3 and 4). The evolution in galaxy structural parameters is rest-frame wavelength-dependent. The redshift and stellar mass dependence of morphological parameters introduce systematic biases in measurements based on single-band photometric images. These biases influence evolutionary trends inferred from scaling relations.
Figure 7: Ratio between sizes measured in the observed $r$-band wavelength and in the rest-frame $r$-band. Points show the median ratios in equally spaced mass bins. Color-coding refers to the four redshift bins in the legend. The shadowed areas indicate the interquartile range of the ratios.
Figure 8: Same as Figure 7, but for Sérsic index.
@@ 5.2.1 Size-stellar mass relation
The size–mass relation is a fundamental ingredient for exploring evolutionary trends in the quiescent galaxy population. Previous investigations demonstrate that the slope of the relation is essentially redshift independent at $z<1.5$ (e.g., Newman et al. 2012; van der Wel et al. 2014; Faisst et al. 2017; Mowla et al. 2019; Barone et al. 2022). As the universe ages, the normalization of the relation increases. The increase reflects the size growth of the population (e.g., Daddi et al. 2005; Trujillo et al. 2007; van Dokkum et al. 2010; Damjanov et al. 2011; Newman et al. 2012; Huertas-Company et al. 2013; van der Wel et al. 2014; Huertas-Company et al. 2015; Faisst et al. 2017; Damjanov et al. 2019; Mowla et al. 2019; Kawinwanichakij et al. 2021; Hamadouche et al. 2022; Damjanov et al. 2023; George et al. 2024; Euclid Collaboration et al. 2025a).
@@
The size-stellar mass relation is a function of galaxy stellar population age (traced by $D_{n}4000$ index). Using quiescent galaxies at $z<0.08$ from SDSS, Zahid & Geller (2017) show that older objects with higher $D_{n}4000$ measurements have smaller sizes. A dense spectroscopic survey spanning the redshift interval $0.50.2<z<$ 0.5 (SHELS F2, Geller et al. 2014, 2016) confirms the anticorrelation between galaxy size (from HSC-SSP $r$-band imaging) and $D_{n}4000$ at intermediate redshift (Damjanov et al. 2018). Damjanov et al. (2023) extends this result to $z\sim 0.6$ using an order of magnitude larger statistical sample of HectoMAP quiescent galaxies with $i$-band imaging. Here we take the advantage of the multi-wavelength HSC-SSP coverage of the HectoMAP survey to explore constraints on the size-stellar mass relation for quiescent galaxies in two different rest-frame wavelengths. We examine four redshift bins covering the $0.60.2<z<$ 0.6 range.
@@
Figure 9: Size-stellar mass relation for quiescent galaxies in two rest-frames. The black points and red crosses show the median galaxy size (log scale) for the central values of equally spaced stellar mass bins in the two rest-frames $3500$ and $7000$  respectively. The panels show four redshift bins with redshift increasing from left to right. The black solid and red dashed lines show the linear fit to the medians, and the thickness of dark gray and red bands around them corresponds to the uncertainties. The shaded bands represent the interquartile ranges of galaxy sizes across stellar mass bins.
@@
Figure 9 shows the size–stellar mass relation for the mass-limited sample of HectoMAP quiescent objects with size estimates at two different rest-frame wavelengths. In each panel, black points and red crosses show the median sizes in 0.25 dex-wide bins of stellar mass at the rest-frame wavelengths of $3500$ (hereafter “UV”) and $7000$ (hereafter “red”), respectively. We fit the relation $\log R_{e,c}^{RF}=R_{0}^{RF}+\gamma\log{(M_{\star}/10^{11}\text{M}_{\odot})}$ to the median values. The black solid and red dashed lines show the results in the UV and red rest-frames, respectively.
@@
The size-stellar mass relation for quiescent galaxies exhibits a pivot point at $M_{\star}\sim(3-5)\times 10^{10}$. The slope of the relation is shallower for masses below the pivot point and steeper for larger masses (e.g., Shen et al. 2003; Williams et al. 2010; Mowla et al. 2019; Kawinwanichakij et al. 2021; Damjanov et al. 2023; George et al. 2024). In the two lowest redshift bins (first two panels from the left), the lower limit of the stellar mass lies below the pivot point mass. For these two redshifts, we fit the median values only for bins with stellar masses exceeding the pivot point mass. We determine the pivot point by using the software package pwlf (Jekel & Venter 2019) to fit a continuous piecewise linear function limited to two segments with different slopes. The crossing point of these two segments defines the pivot point mass. In Figure 9, the lower extreme of each fitted line identifies the pivot point. At $z>0.4$ the lower limit of the sample completeness is $\sim 1$ dex higher than the completeness limit at $z<0.4$. For $z>0.4$ (two rightmost panels of Figure 9), all of the objects in the samples have stellar masses exceeding the pivot point mass.
@@
The slopes of the size - stellar mass relation for HectoMAP galaxies over the the full redshift interval of the survey ($0.60.2<z<$ 0.6) are $\gamma_{\mathrm{UV}}=0.635\pm 0.009$ and $\gamma_{\mathrm{red}}=0.654\pm 0.009$ in the UV and red rest-frame, respectively. Studies based on the Hubble Space Telescope (HST) imaging and spectroscopy (van der Wel et al. 2014; Mowla et al. 2019) cover $0.50.2<z<$ 0.5 with a single redshift bin containing $\sim 100$ galaxies. In both rest-frame regimes, the HectoMAP size-stellar mass relation agrees (within $\lesssim 1.5\,\sigma$) with the size-stellar mass relation slope of $0.75\pm 0.06$ from the HST-based study in the $5000$ rest-frame (van der Wel et al. 2014). The value of $\gamma_{\mathrm{red}}$ is identical to the slope of the size - stellar mass relation for the HectoMAP galaxies at $0.30.2<z<$ 0.3 with size measurements based on a single $i$-band HSC imaging (Damjanov et al. 2023).
@@
The zero point of the size-stellar mass relation $R_{0}^{RF}$, equivalent to the characteristic size of quiescent galaxies at the fiducial stellar mass, is also consistent with previously reported values based on size measurements from either single rest-frame or single observed frame bands. In the redshift range $0.50.2<z<$ 0.5, van der Wel et al. (2014) reports $\log(R_{0}^{RF,5000\leavevmode\nobreak\ \AA}/1\,\mathrm{kpc})=0.60\pm 0.02$ for a typical $5\times 10^{10}$ galaxy. At the same stellar mass, the characteristic sizes we measure are $\log(R_{0}^{\mathrm{UV}}/1\,\mathrm{kpc})=0.558\pm 0.007$ and $\log(R_{0}^{\mathrm{red}}/1\,\mathrm{kpc})=0.50\pm 0.01$ in the UV and red rest-frame, respectively. The characteristic size of HectoMAP galaxies in UV is thus consistent (at $\lesssim 2\sigma$ level) with the HST-based study at the rest-frame $5000$. In the $0.60.2<z<$ 0.6 interval, the characteristic size of a $10^{11}\,M_{\odot}$ HectoMAP galaxy is $\log(R_{0}^{\mathrm{UV}}/1\,\mathrm{kpc})=0.743\pm 0.007$ and $\log(R_{0}^{\mathrm{red}}/1\,\mathrm{kpc})=0.664\pm 0.006$ in the UV and red rest-frame, respectively. The value based on a single ($i$-band) measurement, $\log(R_{0}/1\,\mathrm{kpc})=0.720\pm 0.004$ (Damjanov et al. 2023), lies between the two rest-frame estimates.
@@
The high density of the HectoMAP survey enables segregation of quiescent galaxies into narrow bins of redshift and stellar mass. In combination with the multi-band HSC images of the survey area, the HectoMAP survey enables examination of differences in the redshift evolution of the size-stellar mass relation at two rest-frame wavelengths. In the red rest-frame (i.e., at $7000$, red crosses and dashed lines in Figure 9), the zero point of the size-mass relation increases from $\log(R_{0}^{\mathrm{red}}/1\,\mathrm{kpc})=0.608\pm 0.007$ to $\log(R_{0}^{\mathrm{red}}/1\,\mathrm{kpc})=0.703\pm 0.005$ ($\sim 26\%$) from $z\sim 0.55$ to $z\sim 0.25$. The slope of the size-stellar mass relation in the red rest-frame wavelength shows no significant trend with redshift.
@@
The increase of the zero point of the size-stellar mass relation with increasing redshift quantifies the size growth of quiescent galaxies. The HectoMAP results in the red rest-frame are consistent with previous studies at $z\lesssim 1$ that use size measurements based on imaging in a similar rest-frame (Faisst et al. 2017; Hamadouche et al. 2022; Damjanov et al. 2023). The negligible evolution in the slope of the relation also agrees with previous work (e.g., Newman et al. 2012; van der Wel et al. 2014; Faisst et al. 2017; Mowla et al. 2019; Barone et al. 2022).
@@
In the UV rest-frame ($3500$) we detect no significant redshift evolution either in the zero point or the slope of the size-stellar mass relation. We confirm previous findings based on purely photometric samples with size estimates in the same rest-frame wavelength regime(George et al. 2024).
@@
Figure 10: Redshift evolution of the zero point of the size-stellar mass relation (i.e., $\log(R_{e}/1\,\mathrm{kpc})$ at the reference stellar mass of $10^{11}$ M) as a function of $D_{n}4000$. The top and bottom panels show the results in the UV and red rest-frames, respectively. Black points show the median size of the total sample of galaxies and the sets of colder and warmer points show the zero points for the subpopulations of newcomers and aging galaxies, respectively. The colors of symbols for the two subpopulations correspond to their median $Dn_{4000}$. The error bars denote the 25th-75th percentile range of the distribution. The lines are power law fits to the evolution of the zero point of the correspondingly colored subsample.
@@
Figure 10 illustrates the size evolution of a typical $10^{11}\,M_{\odot}$ HectoMAP galaxies with different average stellar population ages at two rest-frame wavelengths. To derive the characteristic sizes in the first three redshift bins, we interpolate between the median sizes of the two mass bins with mass nearest to $10^{11}$. We compute the characteristic size as the interpolated value at this fiducial mass. The interpolated values are equivalent to the values of zero points $R_{0}$ from the best-fit linear relations in Figure 9. In the highest redshift bin, the characteristic size is the extrapolation of the size - stellar mass relation defined by the two points at $M_{\star}>10^{11}\,M_{\odot}$.
@@
The black circles with error bars in Figure 10 show the redshift evolution in the characteristic size for the full quiescent sample and its scatter in the UV (upper panel) and red (bottom panel). The black lines show the best-fit relation of the form $R_{0}^{RF}(z)=a(1+z)^{b}$. The slope of the relation in the red rest-frame is $-0.98\pm 0.11$, much steeper than the slope in the UV, $-0.13\pm 0.10$. As redshift decreases from $z\sim 0.55$ to $z\sim 0.25$, the ratio between the characteristic sizes in UV and red decreases from $\sim 1.3$ to $\sim 1.1$.
@@
The change in characteristic size with redshift is similar for the newcomers and for the aging population. In the UV rest-frame, the sizes of two subpopulations are essentially constant over the range $0.60.2<z<$ 0.6. For the red rest-frame wavelength, the sizes of newcomers and aging galaxies increase by $\sim 26-30\%$ from the high-redshift to the low-redshift limit. At fixed mass and redshift, newcomers are larger than aging galaxies in both rest-frames. Characteristic sizes for the total sample always lie between the sizes of these two subpopulations. The offset in size between young quiescent galaxies (newcomers) and their older counterparts (aging population) and the size growth in the rest-frame red light agrees with previous studies based on single-band (visible red or near-infrared) imaging (e.g., Zahid & Geller 2017; Hamadouche et al. 2022; Damjanov et al. 2023).
@@
Minor mergers are widely regarded as the main driver of quiescent galaxy size evolution at $z\lesssim 1$ (Naab et al. 2009; Nipoti et al. 2009, 2012; Newman et al. 2012; Faisst et al. 2017; Hamadouche et al. 2022; Damjanov et al. 2024). In this scenario, the absence of size growth in the rest-frame UV for massive ($M_{\star}\sim 10^{11}\,M_{\odot}$) aging galaxies (purple/red colored points in the top panel of Figure 10) may indicate that the star forming material and/or metal-poor stars are mainly accreted at $z\gtrsim 0.5$ where minor mergers with blue satellites are more frequent (e.g., Lotz et al. 2011; Suess et al. 2023). At $z\lesssim 0.55$, accreted material is consumed in residual star formation episodes in the galaxy outskirts. The metal-poor stars accreted at higher redshift contribute to the galaxy halo light. The constant size of massive newcomers in UV light mirrors the lack of substantial size growth in UV for their massive star-forming counterparts (George et al. 2024).
@@
The size increase with decreasing redshift in the red rest-frame (bottom panel of Figure 10) implies that at lower redshift the distribution of the older stellar population (i.e., the bulk of stellar mass) extends to larger radii. This size growth in red light for the aging population suggests an increase in the fraction of low-mass red (i.e., old and/or metal-rich) satellites merging with massive quiescent galaxies at later epochs. Observing the color evolution of low-mass satellites at intermediate redshift is challenging due to their surface brightness (Guo et al. 2011; Nierenberg et al. 2013). With the promised increase in imaging sensitivity and survey area, developing and ongoing missions (e.g., Euclid, Euclid Collaboration et al. 2024; Vera Rubin Observatory’s LSST, Ivezić et al. 2019) may soon provide large enough samples of low surface brightness systems to test this hypothesis.
@@
Residual star formation has a similar effect on the size growth of both newcomers and aging quiescent galaxies in the red rest-frame. This process adds stellar mass to the galaxy outskirts, increasing galaxy size in the red rest-frame. Regions dominated by blue light are at galactocentric distances larger than the regions encompassed by the Hectospec fibers at intermediate redshift ($0.5-2\leavevmode\nobreak\ R_{e}$ in observed $i-$band, Damjanov et al. 2024) and thus they do not affect the global average stellar population age derived in the HectoMAP survey. Spatially resolved spectroscopic observations are critical for measuring stellar population age (and metallicity) in galaxy outskirts. However, these low surface brightness regions are beyond the reach of currently available integral field unit (IFU) observations in the rest-frame UV/visible light even for $z\sim 0$ galaxies (the limits are $<1.5-2\,R_{e}$, Lacerna et al. 2020; Parikh et al. 2021; Lu et al. 2023).

### v2

@@ 5 Structural evolution of quiescent galaxies in two rest-frame wavelengths
Scaling relations for quiescent galaxies constrain the processes that drive their evolution. We first discuss bias that the observations limited to a single photometric band may introduce in the measured structural parameters of quiescent systems at $0.2<z<0.6$ (Sect. 5.1). We then use rest-frame estimates to explore two scaling relations (size-stellar mass and the Sérsic index-stellar mass) in two wavelength ranges (UV and red). We trace the corresponding structural evolution in size and central concentration of quiescent galaxies over the $0.2<z<0.6$ redshift interval (Sec. 5.2).
@@ 5.1 Sérsic profile parameters in observed and rest-frame wavelengths
Sizes decrease with wavelength and Sérsic index increase with wavelength for quiescent galaxies with $0.2<z<0.6$ (Sect. 4.1, Figs. 3 and 4). The evolution in galaxy structural parameters is rest-frame wavelength-dependent. The redshift and stellar mass dependence of morphological parameters introduces systematic biases in measurements based on single-band photometric images. These biases influence evolutionary trends inferred from scaling relations.
Figure 7: Ratio between sizes measured in the observed $r$-band wavelength and in the rest-frame $r$ band. Points show the median ratios in equally spaced mass bins. Color-coding refers to the four redshift bins in the legend. The shadowed areas indicate the interquartile range of the ratios.
Figure 8: Same as Fig. 7, but the for Sérsic index.
@@ 5.2.1 Size-stellar mass relation
The size–mass relation is a fundamental ingredient for exploring evolutionary trends in the quiescent galaxy population. Previous investigations demonstrate that the slope of the relation is essentially redshift independent at $z<1.5$ (e.g., Newman et al. 2012; van der Wel et al. 2014; Faisst et al. 2017; Mowla et al. 2019; Barone et al. 2022). As the universe ages, the normalization of the relation increases. The increase reflects the size growth of the population (e.g., Daddi et al. 2005; Trujillo et al. 2007; van Dokkum et al. 2010; Damjanov et al. 2011; Newman et al. 2012; Huertas-Company et al. 2013; van der Wel et al. 2014; Huertas-Company et al. 2015; Faisst et al. 2017; Damjanov et al. 2019; Mowla et al. 2019; Kawinwanichakij et al. 2021; Hamadouche et al. 2022; Damjanov et al. 2023; George et al. 2024; Euclid Collaboration et al. 2025b).
@@
The size-stellar mass relation is a function of galaxy stellar population age (traced by $D_{n}4000$ index). Using quiescent galaxies at $z<0.08$ from SDSS, Zahid & Geller (2017) show that older objects with higher $D_{n}4000$ measurements have smaller sizes. A dense spectroscopic survey spanning the redshift interval $0.2<z<0.5$ (SHELS F2, Geller et al. 2014, 2016) confirms the anticorrelation between galaxy size (from HSC-SSP $r$-band imaging) and $D_{n}4000$ at intermediate redshift (Damjanov et al. 2018). Damjanov et al. (2023) extends this result to $z\sim 0.6$ using an order of magnitude larger statistical sample of HectoMAP quiescent galaxies with $i$-band imaging. Here we take the advantage of the multiwavelength HSC-SSP coverage of the HectoMAP survey to explore constraints on the size-stellar mass relation for quiescent galaxies in two different rest-frame wavelengths. We examine four redshift bins covering the $0.2<z<0.6$ range.
@@
Figure 9: Size-stellar mass relation for quiescent galaxies in two rest frames. The black points and red crosses show the median galaxy size (log scale) for the central values of equally spaced stellar mass bins in the two rest frames $3500$ Å and $7000$ Å  respectively. The panels show four redshift bins with redshift increasing from left to right. The black solid and red dashed lines show the linear fit to the medians, and the thickness of dark gray and red bands around them corresponds to the uncertainties. The shaded bands represent the interquartile ranges of galaxy sizes across stellar mass bins.
@@
Figure 9 shows the size–stellar mass relation for the mass-limited sample of HectoMAP quiescent objects with size estimates at two different rest-frame wavelengths. In each panel, black points and red crosses show the median sizes in 0.25 dex-wide bins of stellar mass at the rest-frame wavelengths of $3500$ Å (hereafter, “UV”) and $7000$ Å (hereafter, “red”), respectively. We fit the relation $\log R_{e,c}^{RF}=R_{0}^{RF}+\gamma\log{(M_{\star}/10^{11}\text{M}_{\odot})}$ to the median values. The black solid and red dashed lines show the results in the UV and red rest frames, respectively.
@@
The size-stellar mass relation for quiescent galaxies exhibits a pivot point at $M_{\star}\sim(3-5)\times 10^{10}$M. The slope of the relation is shallower for masses below the pivot point and steeper for larger masses (e.g., Shen et al. 2003; Williams et al. 2010; Mowla et al. 2019; Kawinwanichakij et al. 2021; Nedkova et al. 2021; Damjanov et al. 2023; Cutler et al. 2024; George et al. 2024; Hamadouche et al. 2025). In the two lowest redshift bins (first two panels from the left), the lower limit of the stellar mass lies below the pivot point mass. For these two redshifts, we fit the median values only for bins with stellar masses exceeding the pivot point mass. We determine the pivot point by using the software package pwlf (Jekel & Venter 2019) to fit a continuous piecewise linear function limited to two segments with different slopes. The crossing point of these two segments defines the pivot point mass. In Fig. 9, the lower extreme of each fitted line identifies the pivot point. At $z>0.4$ the lower limit of the sample completeness is $\sim 1$ dex higher than the completeness limit at $z<0.4$. For $z>0.4$ (two rightmost panels of Fig. 9), all of the objects in the samples have stellar masses exceeding the pivot point mass.
@@
The slopes of the size - stellar mass relation for HectoMAP galaxies over the full redshift interval of the survey ($0.2<z<0.6$) are $\gamma_{\mathrm{UV}}=0.635\pm 0.009$ and $\gamma_{\mathrm{red}}=0.654\pm 0.009$ in the UV and red rest frame, respectively. Studies based on the Hubble Space Telescope (HST) imaging and spectroscopy (van der Wel et al. 2014; Mowla et al. 2019) cover $0.2<z<0.5$ with a single redshift bin containing $\sim 100$ galaxies. In both rest-frame regimes, the HectoMAP size-stellar mass relation agrees (within $\lesssim 1.5\,\sigma$) with the size-stellar mass relation slope of $0.75\pm 0.06$ from the HST-based study in the $5000$ Å rest-frame (van der Wel et al. 2014). The value of $\gamma_{\mathrm{red}}$ is identical to the slope of the size - stellar mass relation for the HectoMAP galaxies at $0.2<z<0.3$ with size measurements based on a single $i$-band HSC imaging (Damjanov et al. 2023).
@@
The zero point of the size-stellar mass relation $R_{0}^{RF}$, equivalent to the characteristic size of quiescent galaxies at the fiducial stellar mass, is also consistent with previously reported values based on size measurements from either single rest-frame or single observed frame bands. In the redshift range $0.2<z<0.5$, van der Wel et al. (2014) reports $\log(R_{0}^{RF,5000\penalty 10000\ \AA }/1\,\mathrm{kpc})=0.60\pm 0.02$ for a typical $5\times 10^{10}$M galaxy. At the same stellar mass, the characteristic sizes we measure are $\log(R_{0}^{\mathrm{UV}}/1\,\mathrm{kpc})=0.558\pm 0.007$ and $\log(R_{0}^{\mathrm{red}}/1\,\mathrm{kpc})=0.50\pm 0.01$ in the UV and red rest frame, respectively. The characteristic size of HectoMAP galaxies in UV is thus consistent (at $\lesssim 2\sigma$ level) with the HST-based study at the rest-frame $5000$ Å. In the $0.2<z<0.6$ interval, the characteristic size of a $10^{11}\,M_{\odot}$ HectoMAP galaxy is $\log(R_{0}^{\mathrm{UV}}/1\,\mathrm{kpc})=0.743\pm 0.007$ and $\log(R_{0}^{\mathrm{red}}/1\,\mathrm{kpc})=0.664\pm 0.006$ in the UV and red rest frame, respectively. The value based on a single ($i$-band) measurement, $\log(R_{0}/1\,\mathrm{kpc})=0.720\pm 0.004$ (Damjanov et al. 2023), lies between the two rest-frame estimates.
@@
The high density of the HectoMAP survey enables the segregation of quiescent galaxies into narrow bins of redshift and stellar mass. In combination with the multi-band HSC images of the survey area, the HectoMAP survey enables the examination of differences in the redshift evolution of the size-stellar mass relation at two rest-frame wavelengths. In the red rest frame (i.e., at $7000$ Å, red crosses and dashed lines in Fig. 9), the zero point of the size-mass relation increases from $\log(R_{0}^{\mathrm{red}}/1\,\mathrm{kpc})=0.608\pm 0.007$ to $\log(R_{0}^{\mathrm{red}}/1\,\mathrm{kpc})=0.703\pm 0.005$ ($\sim 26\%$) from $z\sim 0.55$ to $z\sim 0.25$. The slope of the size-stellar mass relation in the red rest-frame wavelength shows no significant trend with redshift.
@@
The increase of the zero point of the size-stellar mass relation with increasing redshift quantifies the size growth of quiescent galaxies. The HectoMAP results in the red rest frame are consistent with previous studies at $z\lesssim 1$ that use size measurements based on imaging in a similar rest-frame (Faisst et al. 2017; Hamadouche et al. 2022; Damjanov et al. 2023). The negligible evolution in the slope of the relation also agrees with previous work (e.g., Newman et al. 2012; van der Wel et al. 2014; Faisst et al. 2017; Mowla et al. 2019; Barone et al. 2022).
@@
In the UV rest-frame ($3500$ Å) we detect no significant redshift evolution either in the zero point or the slope of the size-stellar mass relation. We confirm previous findings based on purely photometric samples with size estimates in the same rest-frame wavelength regime(George et al. 2024).
@@
Figure 10: Redshift evolution of the zero point of the size-stellar mass relation (i.e., $\log(R_{e}/1\,\mathrm{kpc})$ at the reference stellar mass of $10^{11}$ M) as a function of $D_{n}4000$. The top and bottom panels show the results in the UV and red rest frames, respectively. Black points show the median size of the total sample of galaxies, and the sets of colder and warmer points show the zero points for the subpopulations of newcomers and aging galaxies, respectively. The colors of symbols for the two subpopulations correspond to their median $D_{n}4000$. The error bars denote the 25th-75th percentile range of the distribution. The dashed blue and dash-dotted orange lines are power law fits to the evolution of the zero point of the subpopulations of newcomers and aging galaxies, respectively.
@@
Figure 10 illustrates the size evolution of a typical $10^{11}\,M_{\odot}$ HectoMAP galaxies with different average stellar population ages at two rest-frame wavelengths. To derive the characteristic sizes in the first three redshift bins, we interpolate between the median sizes of the two mass bins with mass nearest to $10^{11}$M. We compute the characteristic size as the interpolated value at this fiducial mass. The interpolated values are equivalent to the values of zero points $R_{0}$ from the best-fit linear relations in Fig. 9. In the highest redshift bin, the characteristic size is the extrapolation of the size - stellar mass relation defined by the two points at $M_{\star}>10^{11}\,M_{\odot}$.
@@
The black circles with error bars in Fig. 10 show the redshift evolution in the characteristic size for the full quiescent sample and its scatter in the UV (upper panel) and red (bottom panel). The black lines show the best-fit relation of the form $R_{0}^{RF}(z)=a(1+z)^{b}$. The slope of the relation in the red rest frame is $0.98\pm 0.11$, much steeper than the slope in the UV, $0.13\pm 0.10$. As redshift decreases from $z\sim 0.55$ to $z\sim 0.25$, the ratio between the characteristic sizes in UV and red decreases from $\sim 1.3$ to $\sim 1.1$.
@@
The change in characteristic size with redshift is similar for the newcomers and for the aging population. In the UV rest-frame, the sizes of two subpopulations are essentially constant over the range $0.2<z<0.6$. For the red rest frame wavelength, the sizes of newcomers and aging galaxies increase by $\sim 26-30\%$ from the high-redshift to the low-redshift limit. At fixed mass and redshift, newcomers are larger than aging galaxies in both rest frames. Characteristic sizes for the total sample always lie between the sizes of these two subpopulations. The uncertainties in $D_{n}4000$ measurements (Sect. 4.2) have no impact on these results. The offset in size between young quiescent galaxies (newcomers) and their older counterparts (aging population) and the size growth in the rest-frame red light agrees with previous studies based on single-band (visible red or near-infrared) imaging (e.g., Zahid & Geller 2017; Hamadouche et al. 2022; Damjanov et al. 2023).
@@
Minor mergers are widely regarded as the main driver of quiescent galaxy size evolution at $z\lesssim 1$ (Naab et al. 2009; Nipoti et al. 2009, 2012; Newman et al. 2012; Faisst et al. 2017; Hamadouche et al. 2022; Damjanov et al. 2025). In this scenario, the absence of size growth in the rest-frame UV for massive ($M_{\star}\sim 10^{11}\,M_{\odot}$) aging galaxies (purple and red colored points in the top panel of Fig. 10) may indicate that the star forming material and/or metal-poor stars are mainly accreted at $z\gtrsim 0.5$ where minor mergers with blue satellites are more frequent (e.g., Lotz et al. 2011; Suess et al. 2023). At $z\lesssim 0.55$, accreted material is consumed in residual star formation episodes in the galaxy outskirts. The metal-poor stars accreted at higher redshift contribute to the galaxy halo light. The constant size of massive newcomers in UV light mirrors the lack of substantial size growth in UV for their massive star-forming counterparts (George et al. 2024).
@@
The size increase with decreasing redshift in the red rest frame (bottom panel of Fig. 10) implies that at lower redshift the distribution of the older stellar population (i.e., the bulk of stellar mass) extends to larger radii. This size growth in red light for the aging population suggests an increase in the fraction of low-mass red (i.e., old and/or metal-rich) satellites merging with massive quiescent galaxies at later epochs. Observing the color evolution of low-mass satellites at intermediate redshift is challenging due to their surface brightness (Guo et al. 2011; Nierenberg et al. 2013). With the promised increase in imaging sensitivity and survey area, developing and ongoing missions (e.g., Euclid, Euclid Collaboration et al. 2025a; Vera Rubin Observatory’s LSST, Ivezić et al. 2019) may soon provide large enough samples of low surface brightness systems to test this hypothesis.
@@
Residual star formation has a similar effect on the size growth of both newcomers and aging quiescent galaxies in the red rest frame. This process adds stellar mass to the galaxy outskirts, increasing galaxy size in the red rest frame. Regions dominated by blue light are at galactocentric distances larger than the regions encompassed by the Hectospec fibers at intermediate redshift ($0.5-2\penalty 10000\ R_{e}$ in observed $i$ band, Damjanov et al. 2025) and thus they do not affect the global average stellar population age derived in the HectoMAP survey. Spatially resolved spectroscopic observations are critical for measuring stellar population age (and metallicity) in galaxy outskirts. However, these low surface brightness regions are beyond the reach of currently available integral field unit (IFU) observations in the rest-frame UV/visible light even for $z\sim 0$ galaxies (the limits are $1.5-2\,R_{e}$, Lacerna et al. 2020; Parikh et al. 2021; Lu et al. 2023).